I need to solve by completing the square
X^2+2X+48 I believe I would first divide by 2, but not sure what to do after that. Thank you for the help.

Answer by rothauserc(4718)

we are given x^2 +2x + 48 = 0
subtract 48 from both sides of the =
x^2 +2x = -48
complete the square on the left side of = and add the constant square to the right side
x^2 +2x + 1 = -48 +1
(x + 1)^2 = -47
take the square root of both sides of =
x + 1 = sqrt(-47)
x = sqrt(-47) -1